Short sale example Z Corp 100 Shares 5096 | Initial Margin 30% $100 Maintenance Margin Initial Price 1) Stock Price Rises to $110 2) How much can the stock price rise before a margin call?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins as Cut aCopy Σ AutoSum ー E ゴWrap Text B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Alignment Number Cells Edting 268 268 269 270 271 272 273 274 275 276 277 278 279 280 281 282 283 284 285 286 4 EV PE EPS DIV ABNORMAL APV ROSS deberk MARGIN MONEY PROCEEDS FROM SHORT SALE = 100 X 100 INITIAL MARGIN-50% INTIAL MARGIN DEPOSIT OWN FUNDS 10000 5000 TOTAL ASSETS TOTAL LIABILITIES + EQUITY 10000 + 5000 = 15000 PRICE RISES TO 110 COST OF COVERING SHORT-100x110 11000 4000 REMAINING MARGIN IN ACCOUNT REMAINING MARGIN = ACCOUNT EQUITY = TOTAL ASSETS-COST OF COVERING SHORT ACCOUNT EQUITY 15000 11000 LOSSCOST OF COVERING SHORT - PROCEEDS FROM SHORT SALES LOSS = 11000-10000 = 1000 RATE OF RETURN-PROFIT OR LOSS/OWN FUNDS=-1000/5000 = 20.00% ANSWER -20.00% al tax . PB ROPİ eEVA MVA LEVERAGED BUY SWAP distress BRIGHAN CAP GAINm 03:11 28-01-2019Home nert Page Layout Formulas Data Review View dd-Ins s Cut aCopy E AutoSum ー E ゴWrap Text ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B 1 u. Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Alignment Number Cells Edting K305 287 288 289 290 291 292 293 294 295 296 297 298 299 300 301 302 303 304 305 4 EV PE EPS DIV ABNORMAL APV ROSS deberk MARGIN MONEY MARGIN CALL WILL BE AT A PRICE = (TOTAL ASSETS /NO OF SHARES)/(1 + MAINTENANCE MARGIN) MARGIN CALL WILL BE AT A PRICE-(15000/100)/(1+0.30)- 115.38 al tax . PB ROPİ eEVA MVA LEVERAGED BUY SWAP distress BRIGHAN CAP GAINm 福 130% 03:12